# Infinite-Dimensional Stochastic Differential Equations and Diffusion Dynamics of Coulomb Random Point Fields

### v3

@@ 4. Construction of the logarithmic derivative $\mathsf{d}^{\mu}$ of $\mu$
In this section, we establish the existence of the logarithmic derivative of $\mu$. The section is divided into four parts.
In Subsection 4.1, we prove the $L^{2}$-uniform bound of the logarithmic derivatives $\mathsf{d}^{N}$ of the $N$-particle systems. In Subsection 4.2, we demonstrate the $L^{2}$-uniform bound of the logarithmic derivatives $\mu_{R}^{N}$ for $N,R\in\mathbb{N}$. In Subsection 4.3, we show the strong convergence of the density functions of $\mu_{R}^{N}$ as $N\to\infty$ for each fixed $R\in\mathbb{N}$. Finally, in Subsection 4.4, we construct the logarithmic derivative $\mathsf{d}^{\mu}$ of $\mu$.
@@ 4.1. $L^{2}$-uniform bound of the logarithmic derivative $\mathsf{d}^{N}$ of $\mu^{N}$
Let $\mu^{N}$ be a random point field satisfying ((A1))((A3)). Let $\mathfrak{m}^{N}$ be the labeled density function of $\mu^{N}$ given by (2.18). For ${x}=({x}_{1},\ldots,{x}_{d})\in\mathbb{R}^{d}$ and $\mathsf{s}=\sum_{i=1}^{N-1}\delta_{s^{i}}$, we set
$\mathsf{d}^{N}({x},\mathsf{s})=\nabla_{{x}}\log\mathfrak{m}^{N}({x},s^{1},\ldots,s^{N-1}),$ (4.1)
where $\nabla_{{x}}=(\frac{\partial}{\partial{x}_{1}},\ldots,\frac{\partial}{\partial{x}_{d}})$. From (2.18) and (4.1), we have
$\mathsf{d}^{N}({x},\mathsf{s})=-\beta\Big{(}\nabla\Phi^{N}({x})+\sum_{i=1}^{N-1}\nabla_{{x}}\Psi^{N}({x},s^{i})\Big{)}.$ (4.2)
Integration by parts yields, for any $\mathsf{h}\in C_{0}^{\infty}(O_{N})\otimes\mathscr{D}_{\bullet\mathrm{b}}$,
$\int_{\mathbb{R}^{d}\times\mathsf{S}}\mathsf{h}({x},\mathsf{s})\mathsf{d}^{N}({x},\mathsf{s})\mu^{N,[1]}(dxd\mathsf{s})=-\int_{\mathbb{R}^{d}\times\mathsf{S}}\nabla_{{x}}\mathsf{h}({x},\mathsf{s})\mu^{N,[1]}(dxd\mathsf{s}).$
Here, $\mu^{N,[1]}$ is the one-reduced Campbell measure of $\mu$, defined as $d\mu^{N,[1]}=\rho^{N,1}(x)\mu^{N}_{x}(d\mathsf{s})d{x}$, where $\rho^{N,1}$ is the one-point correlation function of $\mu^{N}$ and $\mu^{N}_{x}=\mu^{N}(\cdot+\delta_{{x}}|\mathsf{s}(\{{x}\}\geq 1))$ is the reduced Palm measure conditioned at x{x}. Thus, $\mathsf{d}^{N}$ is the $\bullet$-logarithmic derivative of $\mu$ in the sense of Definition 2.2.
@@
Assume ((A1))((A3)). Let $\mathsf{d}^{N}=(\mathsf{d}_{p}^{N})_{p=1}^{d}$ and $\partial_{p}=\partial/\partial x_{p}$ for ${x}=(x_{p})\in\mathbb{R}^{d}$. Then, for any $\mathsf{h}\in C_{0}^{\infty}(O_{N})\otimes\mathscr{D}_{\bullet\mathrm{b}}$,
$-\int_{\mathbb{R}^{d}\times\mathsf{S}}\mathsf{h}\sum_{p=1}^{d}\partial_{p}\mathsf{d}_{p}^{N}d\mu^{N,[1]}=\beta\int_{\mathbb{R}^{d}\times\mathsf{S}}\mathsf{h}\big{(}\Delta\Phi^{N}+\sum_{i=1}^{N-1}\psi^{N}({x},s^{i})\big{)}d\mu^{N,[1]}.$ (4.3)
@@
$\beta\int_{\mathbb{R}^{d}\times\mathsf{S}}\mathsf{h}({x},\mathsf{s})\sum_{i=1}^{N-1}\Delta_{x}\Psi^{N}({x},s^{i})d\mu^{N,[1]}=\beta\int_{\mathbb{R}^{d}\times\mathsf{S}}\mathsf{h}({x},\mathsf{s})\sum_{i=1}^{N-1}\psi^{N}({x},s^{i})d\mu^{N,[1]}.$ (4.4)
Then, (4.3) follows immediately from (4.2) and (4.4). ∎

### v4

@@ 4. Construction of the logarithmic derivative $\mathsf{d}^{\mu}$ of $\mu$
In this section we establish the existence of the logarithmic derivative $\mathsf{d}^{\mu}$ of $\mu$. The section is organized into four parts.
In Subsection 4.1, we prove a uniform $L^{2}$-bound for the logarithmic derivatives $\mathsf{d}^{N}$ of the $N$-particle systems. In Subsection 4.2, we derive a uniform $L^{2}$-bound for the logarithmic derivatives of $\mu_{R}^{N}$. In Subsection 4.3, we prove the strong convergence of the density functions of $\mu_{R}^{N}$ as $N\to\infty$ for each fixed $R\in\mathbb{N}$. Finally, in Subsection 4.4, we construct the logarithmic derivative $\mathsf{d}^{\mu}$ of $\mu$.
All results in Section 4 are proved under Assumptions ((A1))((A3)).
@@ 4.1. $L^{2}$-uniform bound of the logarithmic derivative $\mathsf{d}^{N}$ of $\mu^{N}$
Let $\mu^{N}$ be as in ((A1))((A3)). Let $\mu^{N,[1]}$ denote the one-reduced Campbell measure associated with $\mu^{N}$, i.e., the specialization of the definition (2.1) (given for a general RPF $\nu$) to the case $\nu=\mu^{N}$.
Let $\mathfrak{m}^{N}$ be the labeled density function of $\mu^{N}$ given by (1.3). Let
$\mathsf{d}^{N}({x},\mathsf{s})=\begin{cases}\nabla_{{x}}\log\mathfrak{m}^{N}({x},s^{1},\ldots,s^{N-1}),&{x}\in O_{N},\ \mathsf{s}(\mathbb{R}^{d})=N-1,\\ 0,&\text{otherwise},\end{cases}$ (4.1)
where $\mathsf{s}=\sum_{i=1}^{N-1}\delta_{s^{i}}$. By (1.3) and (4.1),
$\mathsf{d}^{N}({x},\mathsf{s})=-\beta\Big(\nabla\Phi^{N}({x})+\sum_{i=1}^{N-1}\nabla_{{x}}\Psi^{N}({x},s^{i})\Big),\quad{x}\in O_{N}.$ (4.2)
@@
For any $\mathsf{h}\in C_{0}^{\infty}(O_{N})\otimes\mathscr{D}_{\bullet\mathrm{b}}$,
$-\int_{\mathbb{R}^{d}\times\mathsf{S}}\mathsf{h}\,(\mathrm{div}_{x}\mathsf{d}^{N})\,d\mu^{N,[1]}=\beta\int_{\mathbb{R}^{d}\times\mathsf{S}}\mathsf{h}\Big(\Delta\Phi^{N}+\sum_{i=1}^{N-1}\psi^{N}({x},s^{i})\Big)d\mu^{N,[1]}.$
@@
Moreover, by ((A1)) and ((A2)), the density $\mathfrak{m}^{N}$ vanishes on collision configurations, that is, $\mathfrak{m}^{N}(\mathbf{x})=0$ whenever $x^{i}=x^{j}$ for some $i\neq j$. Hence the singular delta terms in (4.3) do not contribute, and we obtain
$\int_{\mathbb{R}^{d}\times\mathsf{S}}\mathsf{h}\sum_{i=1}^{N-1}\Delta_{{x}}\Psi^{N}({x},s^{i})\,d\mu^{N,[1]}=\int_{\mathbb{R}^{d}\times\mathsf{S}}\mathsf{h}\sum_{i=1}^{N-1}\psi^{N}({x},s^{i})\,d\mu^{N,[1]}.$
Combining this with (4.2), the assertion follows. ∎